# Gradient Descent with Large Step Sizes: Chaos and Fractal Convergence Region

### v1

@@ E.2.1 Preliminary results
which clearly holds when $z$ is sufficiently small. Therefore, $m_{k}$ must lie in the region $\left\{w^{\prime}>w\right\}$ for all $k\geq K$ for some finite $K>0$. Note this implies that the $w$-coordinate of $m_{k}$ starts to decrease from all $k\geq K$. This conflicts with the fact that $m_{k}$ is unbounded. Hence $m_{k}$ is bounded.
According to (19), $m_{k}$ has to converge to
$\mathrm{cl}(A_{0})\cap\{(w+2(z+\mu))(z+\nu-2)(z+\nu)=0\}=\{(w+2(z+\mu)=0\}.$
Otherwise, assume $m_{k}\subset K$ for all $k$ and some compact set $K$. The function E(F(z,w))E(z,w)E(F(z,w))-E(z,w) is uniform continuous, so if $m_{k}$ does not converge to its zero set, $E(m_{k})-E(m_{k-1})$ is bounded below and $E(m_{k})$ can not converge.
Note, when restricting to $w=-2(z+\mu)$, the $w$-update is given by
$w^{\prime}=d(w)=w(\frac{w}{2}-1+\nu+\mu)^{2}.$
Solving $d(w)=w$, we obtain that $d$ has three fixed points, $0,-2(\mu+\nu)$ and $4-2(\mu+\nu)$. Note when $\mu\geq 0$, we have $2(\mu+\nu)<0$. Therefore, on $w\geq 0$, $d$ has two fixed points. Note $d^{\prime}(0)=(-1+\mu+\nu)^{2}<1$ as $\mu+\nu<1$, and hence $w=0$ is repelling under $d^{-1}$. Note $d^{\prime}(4-2(\mu+\nu))=5-2(\mu+\nu)>1$, and hence $w=4-2(\mu+\nu)$ is attracting under $d^{-1}$. With basic graph analyses, we have that $w=4-2(\mu+\nu)$ globally attracts all orbits under $d^{-1}$, except the point $w=0$. Notice that, one must have $m_{1}\neq(-\mu,0)$ as $m_{1}\in\mathrm{cl}(A_{1})$. Therefore, since $m_{k}$ converges to $w=-2(z+\mu)$, we have that $m_{k}$ converges to $(-2+\nu,4-2(\nu+\mu))$. The case of $y<0$ can be proved via a analogous procedure. This completes the proof. ∎
@@
Consider $\xi_{1}=(-2+\nu,4-2(\nu+\mu))$ and $\xi_{2}=(2-\nu,4+2(\nu+\nu))$. For $i=1,2$, we have that $\cup_{N=0}^{\infty}F^{-N}(\xi_{i})$ has infinitely many points and $\cup_{N=0}^{\infty}F^{-N}(\xi_{i})\subset\partial T(\mathcal{D}_{\eta}^{\prime\prime})$. When $y\geq 0$, for any open set $O$ such that $O\cap\cup_{N=0}^{\infty}F^{-N}(\xi_{1})\neq\varnothing$, there exists $(z^{\prime},w^{\prime}),(z^{\prime\prime},w^{\prime\prime})\in O$ such that, for any $(u^{\prime},v^{\prime}),(u^{\prime\prime},v^{\prime\prime})$ that satisfy $T(u^{\prime},v^{\prime})=(z^{\prime},w^{\prime})$ and $T(u^{\prime\prime},v^{\prime\prime})=(z^{\prime\prime},w^{\prime\prime})$, we have $\mathrm{GD}_{\eta}^{N}(u^{\prime},v^{\prime})$ converges $p^{+}(u^{\prime},v^{\prime})$ and $\mathrm{GD}_{\eta}^{N}(u^{\prime\prime},v^{\prime\prime})$ converges $p^{-}(u^{\prime\prime},v^{\prime\prime})$. When $y<0$, the same result holds for any open set $O$ such that $O\cap\cup_{N=0}^{\infty}F^{-N}(\xi_{2})\neq\varnothing$.
@@
We present the proof for $y\geq 0$ for brevity. The case $y<0$ can be proved via an analogous procedure. We first show that $\cup_{N=0}^{\infty}F^{-N}(\xi_{i})$ has infinitely many points and $\cup_{N=0}^{\infty}F^{-N}(\xi_{i})\subset\partial T(\mathcal{D}_{\eta}^{\prime\prime})$. Notice that, $\xi_{1}$ lies in the set $\left\{(z,w)\in\Omega\colon w=-2(z+\mu)\right\}$. By Proposition 11, this set is invariant under $F$, where the $w$-update is given by
$w^{\prime}=w(\frac{w}{2}-1+\mu+\nu)^{2}.$
By differentiation, we know $w=4-2(\mu+\nu)$ is a repelling fixed point. Specifically, we claim that on any neighborhood of $4-2(\mu+\nu)$ there exists a point that converges to 0 and a point that diverges to infinity. Note $w$ converging to zero corresponds to $(z,w)$ converging to $(-\mu,0)$, and to $(u,v)$ converging to the saddle $(\boldsymbol{0},\boldsymbol{0})$. Therefore, we have $\xi_{1}\in\partial T(\mathcal{D}_{\eta}^{\prime\prime})$. Note, $Q(\xi_{1})=8-4\nu>6-4\nu$. Therefore, by Proposition 13, $F^{-1}(\xi_{1})$ can be explicitly given by $\cup_{k=0,1,2}G_{i}(\xi_{1})$, where $G_{i}$ is a homeomorphism for all $i$. Therefore,
$\cup_{N=0}^{\infty}F^{-N}(\xi_{1})=\left\{G_{i_{1}}\circ G_{i_{k}}(\xi_{1})\colon\forall k\geq 1,i_{j}\in\left\{0,1,2\right\},\forall j\right\}.$ (20)
By the construction of $G_{i}$, the cardinality of this set is infinity. Also, as each $G_{i}$ is a homeomorphism, any point in this set belongs to $\partial T(\mathcal{D}_{\eta}^{\prime\prime})$.
Next, we show that for any open set $O$ such that $O\cap\cup_{N=0}^{\infty}F^{-N}(\xi_{1})\neq\varnothing$, there exists $(z^{\prime},w^{\prime}),(z^{\prime\prime},w^{\prime\prime})\in O$ satisfy the claimed properties. Note, by (20), it suffices to prove this result for $O$ that satisfies $O\ni\xi_{1}$. When $y\leq\lambda$, $L$ has a unique minimizer $(\boldsymbol{0},\boldsymbol{0})$ and the result obvious holds from Proposition 25. Now consider $y>\lambda$, in which case the set of global minimizers is given by $\left\{q=0,\|p\|^{2}=2(y-\lambda)\right\}$. Consider $p=(u+v)/\sqrt{2}$ and $q=(u-v)/\sqrt{2}$. We have that
$p_{k}=p_{0}\Pi_{j=0}^{k-1}(1-z_{j}-\nu).$
Therefore, the orientation of $p_{\infty}$ is fully determined by whether number of $z_{j}$ such that $1-z_{j}-\nu>0$ is even or odd. Now let $m_{*}=(-\nu,2(\mu-\nu))$ denote the minimizer. Note, the boundary $w=2(z+\mu)$ is invariant by Lemma 11, where the $w$-update is given by
@@
Note these roots are reals since $\mu-\nu>0$. Note on the line $w=2(z+\mu)$, $1-\nu=z$ yields $w=2(1-\nu+\mu)$. Let $\alpha=\mu-\nu$. We have that
@@
which is true since $0<\nu,\mu<1$. Similarly, we have
@@
which is true since $\alpha>0$. Therefore, there exists $m_{*}^{+},m_{*}^{-}\in\Omega$ such that, $F(m_{*}^{\pm})=m^{*}$ and $m_{*}^{+}\in\left\{z>1-\nu\right\}$ and $m_{*}^{-}\in\left\{z,1-\nu\right\}$. By Proposition 25, $G_{0}^{N}(m_{*}^{\pm})$ converges to $(2-\nu,4+2(\nu+\mu))$. Notice that, for $N\geq 1$, $G_{0}^{N}(m_{*}^{\pm})\in\left\{z<1-\nu\right\}$ since the image set of $G_{0}$ is $\mathrm{cl}(A_{0})$. Therefore, the entire sequence $G_{0}^{N}(m_{*}^{+})$ enters $\{z>1-\nu\}$ exactly one time. Similarly, the entire sequence $G_{0}^{N}(m_{*}^{-})$ enters $\{z>1-\nu\}$ exactly zero time. Therefore, the corresponding $uv$-orbit converges the maximal and minimal distance solution. This completes the proof. ∎
E.2.2 Proof of Theorem 3
@@
In the quotient system $F$, the basin of attraction of the point $(-\mu,0)$ has measure zero. This is because as in the proof for Proposition 22, the basin can be given by $\cup_{N=0}^{\infty}F^{-N}(O\cap\left\{w=-2\mathrm{sgn}(\mu)(z+\mu)\right\})$ for some neighborhood $O$ of $(-\mu,0)$. Since the Jacobian of $F$ has full rank almost everywhere, this basin of attraction is a measure zero set. According to Proposition 6, any measure-zero event in system $F$ corresponds to a measure-zero even in system $\mathrm{GD}_{\eta}$. Therefore, $\mathcal{S}_{\eta}$ has measure zero. The projected boundary is self similar with degree three is directly given by Proposition 23. The unboundedness is given by Proposition 24.
@@
Finally, we prove sensitivity to initialization. Consider any open neighborhood $W\subset\mathbb{R}^{2d}$ such that $W\cap\partial\mathcal{D}^{\prime\prime}_{\eta}\neq\varnothing$. We claim that $T(W)$ is a neighborhood that contains an open neighborhood on $T(\partial\mathcal{D}^{\prime\prime}_{\eta})$. Notice the Jacobian of the map $T$ drops rank if and only if $u=\pm v$. If $W\cap\left\{u=\pm v\right\}=\varnothing$, then by constant rank theorem, $T$ is locally a projection, which gives the claim. If $W\cap\left\{u=\pm v\right\}\neq\varnothing$, then without loss of generality, assume $W=B((u_{0},u_{0}),\delta)$. Then $T(u_{0},u_{0})=(\|u_{0}\|^{2},2\|u_{0}\|^{2})$. We show that for any point $(z^{\prime},w^{\prime})$ that is sufficiently close to $(\|u_{0}\|^{2},2\|u_{0}\|^{2})$, there exists a preimage under $T$ in $W$. Note, as $T$ is surjective, assume $T(u^{\prime},v^{\prime})=(z^{\prime},w^{\prime})$. Note whenever $(z^{\prime},w^{\prime})$ tends to $(\|u_{0}\|^{2},2\|u_{0}\|^{2})$, we have $w^{\prime}+2z^{\prime}=\|u^{\prime}+v^{\prime}\|^{2}$ tends to $4\|u_{0}\|^{2}$ and $w^{\prime}-2z^{\prime}=\|u^{\prime}-v^{\prime}\|^{2}$ tends to 0. Therefore, $(u^{\prime},v^{\prime})$ tends to $\left\{u=v,\|u\|=\|u_{0}\|\right\}$. Note, the map $T$ is invariant under rotation. Therefore, with proper rotation, $(u^{\prime},v^{\prime})$ tends to $\left\{u=v,u=u_{0}\right\}=(u_{0},u_{0})$ and thus it lies in $W$. When $y\geq 0$, consider the set $H=T^{-1}(\cup_{N=0}^{\infty}F^{-N}(\xi_{1}))$, where $\xi_{1}$ is defined in Proposition 26. By Proposition 26, $H$ has infinitely many elements. As shown in Proposition 23, we have $T(\partial\mathcal{D}_{\eta}^{\prime\prime})=\partial T(\mathcal{D}_{\eta}^{\prime\prime})$. Hence, $\cup_{N=0}^{\infty}F^{-N}(\xi_{1})\subset\partial T(\mathcal{D}_{\eta}^{\prime\prime})=T(\partial\mathcal{D}_{\eta}^{\prime\prime})$. Therefore, when $W\cap H\neq\varnothing$, $T(W)$ is a neighborhood that intersects with $\cup_{N=0}^{\infty}F^{-N}(\xi_{1})$. By Proposition 26, there exist $\theta^{\prime},\theta^{\prime\prime}\in W$ such that $\mathrm{GD}_{\eta}^{N}(\theta^{\prime})$ converges to $p^{+}(\theta^{\prime})$ and $\mathrm{GD}_{\eta}^{N}(\theta^{\prime\prime})$ converges to $p^{-}(\theta^{\prime})$. The case of $y<0$ can be proved analogously using Proposition 26. This completes the proof. ∎

### v2

@@ E.2.1 Preliminary results
which clearly holds as $z$ tends to -$\infty$. Therefore, $m_{k}$ must lie in the region $\left\{w^{\prime}>w\right\}$ for all $k\geq K$ for some finite $K>0$. Note this implies that the $w$-coordinate of $m_{k}$ starts to decrease from all $k\geq K$. This conflicts the fact that $m_{k}$ is unbounded, as $\Omega\cap\left\{w<C\right\}$ is bounded for any $C>0$. Hence, $m_{k}$ is bounded.
According to (24), $m_{k}$ must converge to
$\mathrm{cl}(A_{0})\cap\{(w+2(z+\mu))(z+\nu-2)(z+\nu)=0\}=\mathrm{cl}(A_{0})\cap\{(w+2(z+\mu)=0\}.$
Otherwise, assume $m_{k}\subset K$ for all $k$ and some compact set $K$. The function E(F(z,w))E(z,w)E(F(z,w))-E(z,w) is continuous, so if $m_{k}$ does not converge to its zero level set, $E(m_{k})-E(m_{k-1})$ is bounded below and $E(m_{k})$ can not converge.
Note, when restricting to $\{w=-2(z+\mu)\}$, the $w$-update under $F$ is given by
$w^{\prime}=\kappa(w)=w(\frac{w}{2}-1+\nu+\mu)^{2},\ w\geq 0.$
Solving $\kappa(w)=w$, we obtain that $\kappa$ has two fixed points on $w\geq 0$: $w=0$ and $w=4-2(\mu+\nu)$. It is straight forward to obtain that all backward orbits of $\kappa$ converges to $w=4-2(\mu+\nu)$ except the one initialized at $w=0$. Meanwhile, note that $w=4-2(\mu+\nu)$ corresponds to $(\nu-2,4-2(\mu+\nu))$. The eigenvalues of the Jacobian of $F$ at this point are: $5-2\mu-2\nu,(-3+2\nu)^{2}$. Therefore, the backward orbits of $F$, i.e., forward orbits of $G_{0}$, are locally attracted by $(\nu-2,4-2(\mu+\nu))$. Then, using omega-limit set arguments similar to those in the proof of Proposition 15, we have that $m_{k}$ converges to $(-2+\nu,4-2(\nu+\mu))$. This completes the proof. ∎
@@
Assume $0\leq\nu<\min\{\frac{1}{2},1-|\mu|\}$. Consider $\xi_{1}=(-2+\nu,4-2(\nu+\mu))$ and $\xi_{2}=(2-\nu,4+2(\nu+\nu))$. For $i=1,2$, we have that $\cup_{N=0}^{\infty}F^{-N}(\xi_{i})$ has infinitely many points and $\cup_{N=0}^{\infty}F^{-N}(\xi_{i})\subset T(\partial\mathcal{D}_{\eta}^{\prime\prime})$. When $y\geq 0$, for any open set $O$ such that $O\cap(\cup_{N=0}^{\infty}F^{-N}(\xi_{1}))\neq\varnothing$, there exists $(z^{\prime},w^{\prime}),(z^{\prime\prime},w^{\prime\prime})\in O$ such that, for any $(u^{\prime},v^{\prime}),(u^{\prime\prime},v^{\prime\prime})$ that satisfy $T(u^{\prime},v^{\prime})=(z^{\prime},w^{\prime})$ and $T(u^{\prime\prime},v^{\prime\prime})=(z^{\prime\prime},w^{\prime\prime})$, we have $\mathrm{GD}_{\eta}^{N}(u^{\prime},v^{\prime})$ converges to $p^{+}(u^{\prime},v^{\prime})$ and $\mathrm{GD}_{\eta}^{N}(u^{\prime\prime},v^{\prime\prime})$ converges to $p^{-}(u^{\prime\prime},v^{\prime\prime})$. When $y<0$, the same result holds for any open set $O$ such that $O\cap(\cup_{N=0}^{\infty}F^{-N}(\xi_{2}))\neq\varnothing$.
@@
We prove the case $y\geq 0$. Note that the case of $y<0$ can be proved via analogous procedures. We first show that $\cup_{N=0}^{\infty}F^{-N}(\xi_{i})$ has infinitely many points and $\cup_{N=0}^{\infty}F^{-N}(\xi_{i})\subset\partial T(\mathcal{D}_{\eta}^{\prime\prime})$. Notice that, $\xi_{1}$ lies in the set $\left\{(z,w)\in\Omega\colon w=-2(z+\mu)\right\}$. By Proposition 11, this set is invariant under $F$, where the $w$-update under $F$ is given by
$w^{\prime}=\kappa(w)=w(\frac{w}{2}-1+\mu+\nu)^{2}.$
By analyzing the one-dimensional cubic map $\kappa$, it is straight forward to obtain that, as $N\to\infty$, for all $w$ with $w<4-2(\mu+\nu)$, $\kappa^{N}(w)\to 0$ and, for all $w$ with $w>4-2(\mu+\nu)$, $\kappa^{N}(w)\to+\infty$. Note that $w$ converging to zero corresponds to $zw$-orbit converging to $(-\mu,0)$ and $uv$-orbit converging to the $(\boldsymbol{0},\boldsymbol{0})$. Therefore, we have $\xi_{1}\in\partial T(\mathcal{D}_{\eta}^{\prime\prime})=T(\partial\mathcal{D}_{\eta}^{\prime\prime})$, where the equality was shown in the proof of Proposition 23. Note, $Q(\xi_{1})=8-4\nu>6-4\nu$. Therefore, by Proposition 13, Therefore,
$\cup_{N=0}^{\infty}F^{-N}(\xi_{1})=\left\{G_{i_{1}}\circ\cdots\circ G_{i_{k}}(\xi_{1})\colon\forall k\geq 1,i_{j}\in\left\{0,1,2\right\},\forall j\right\},$ (25)
where $G_{i}$’s are homeomorphisms. By the construction of $G_{i}$, the cardinality of this set is infinity. Also, as each $G_{i}$ is a homeomorphism, any point in this set belongs to $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$.
Next, we show that for any open set $O$ such that $O\cap\cup_{N=0}^{\infty}F^{-N}(\xi_{1})\neq\varnothing$, there exists $(z^{\prime},w^{\prime}),(z^{\prime\prime},w^{\prime\prime})\in O$ satisfying the claimed properties. Note, by (25), it suffices to prove this for $O$ that satisfies $O\ni\xi_{1}$. When $y\leq\lambda$, $L$ has a unique minimizer $(\boldsymbol{0},\boldsymbol{0})$ and the result obviously holds from Proposition 25. Now consider $y>\lambda$. Let $(u_{k},v_{k})_{k\geq 0}$ be a gradient descent orbit that converges to a global minimizer, and $(z_{k},w_{k})=T(u_{k},v_{k})$. As we shown in the proof of Proposition 22, the converged minimizer is the minimal distance solution if
$\left\{j\geq 0\colon 1-z_{j}-\nu<0\right\}$
is an even number; and the converged minimizer is the maximal distance solution if the above is an odd number. Now let $m_{*}=(-\nu,2(\mu-\nu))$, where $\{m^{*}\}=T(\mathcal{M})$. Note, the boundary $\{w=2(z+\mu)\}$ is invariant under $F$ by Lemma 11, where the $w$-update under $F$ is given by
@@
Note these roots are reals since $\mu-\nu>0$. Note on the line $\{w=2(z+\mu)\}$, $1-\nu=z$ yields $w=2(1-\nu+\mu)$. Let $\alpha=\mu-\nu$. We have that
@@
which is true by assumption. Similarly, we have
@@
which is also true as $\alpha>0$. Therefore, there exists $m_{*}^{+},m_{*}^{-}\in\Omega$ such that, $F(m_{*}^{\pm})=m^{*}$ and $m_{*}^{+}\in\left\{z>1-\nu\right\}$ and $m_{*}^{-}\in\left\{z<1-\nu\right\}$. By Proposition 25, $G_{0}^{N}(m_{*}^{\pm})$ converges to $(2-\nu,4+2(\nu+\mu))$ as $N\to\infty$. This means that we can find $(z^{\prime},w^{\prime})$ and $(z^{\prime\prime},w^{\prime\prime})\in O$ such that $(z^{\prime},w^{\prime})=G_{0}^{N^{\prime}}(m_{*}^{+})$ and $(z^{\prime\prime},w^{\prime\prime})=G_{0}^{N^{\prime\prime}}(m_{*}^{-})$ for some $N^{\prime},N^{\prime\prime}>0$. Notice that, for $N\geq 1$, $G_{0}^{N}(m_{*}^{\pm})\in\left\{z<1-\nu\right\}$ since the image set of $G_{0}$ is $\mathrm{cl}(A_{0})$. Therefore, the $F$-orbit initialized at $(z^{\prime},w^{\prime})$ enters the region $\{z>1-\nu\}$ exactly once, while the $F$-orbit initialized at $(z^{\prime\prime},w^{\prime\prime})$ never enters $\{z>1-\nu\}$. Therefore, the corresponding $uv$-orbit converges to the maximal and minimal distance solutions, respectively. This completes the proof. ∎
E.2.2 Proofs of Theorem 3 and Theorem 4
@@
By Proposition 8, $\mathcal{S}_{\eta}=T^{-1}(T(\mathcal{S}_{\eta}))$, and $T(\mathcal{S}_{\eta})$ is the basin of attraction of the point $(-\mu,0)$ for system $F$. As shown in the proof of Proposition 22, $T(\mathcal{S}_{\eta})$ has measure zero. By Proposition 6, $\mathcal{S}_{\eta}$ also has measure zero. The projected boundary $T(\mathcal{D}_{\eta}^{\prime\prime})$ is self similar with degree three by Proposition 23. The unboundedness is given by Proposition 24.
When $y\geq 0$, consider the set $H=T^{-1}(\cup_{N=0}^{\infty}F^{-N}(\xi_{1}))$, where $\xi_{1}$ is defined in Proposition 26. By Proposition 26 and since $T$ is surjective, $H$ has infinitely many elements. By Proposition 8, $T^{-1}(T(\partial\mathcal{D}_{\eta}^{\prime\prime}))=\partial\mathcal{D}_{\eta}^{\prime\prime}$. By Proposition 26, $\cup_{N=0}^{\infty}F^{-N}(\xi_{1})\subset T(\partial\mathcal{D}_{\eta}^{\prime\prime})$. Together, we have that $H=T^{-1}(\cup_{N=0}^{\infty}F^{-N}(\xi_{1}))\subset\partial\mathcal{D}_{\eta}^{\prime\prime}$.
@@
Consider any open neighborhood $W\subset\mathbb{R}^{2d}$ such that $W\cap H\neq\varnothing$. We will show that $T(W)$ contains an open neighborhood $O$ such that $O\cap(\cup_{N=0}^{\infty}F^{-N}(\xi_{1}))\neq\varnothing$. Notice the Jacobian of the map $T$ drops rank if and only if $u=\pm v$. If $W\cap\left\{u=\pm v\right\}=\varnothing$, then by constant rank theorem, $T$ is locally a projection, which gives the claim. If $W\cap\left\{u=\pm v\right\}\neq\varnothing$, then, without loss of generality, assume $W=B((u_{0},u_{0}),\delta)$. Then $T(u_{0},u_{0})=(\|u_{0}\|^{2},2\|u_{0}\|^{2})$. We show that for any point $(z^{\prime},w^{\prime})\in\Omega$ that is sufficiently close to $(\|u_{0}\|^{2},2\|u_{0}\|^{2})$, there exists a preimage under $T$ in $W$. Note, as $T$ is surjective, there exists $(u^{\prime},v^{\prime})$ such that $T(u^{\prime},v^{\prime})=(z^{\prime},w^{\prime})$. Note whenever $(z^{\prime},w^{\prime})$ tends to $(\|u_{0}\|^{2},2\|u_{0}\|^{2})$, we have $w^{\prime}+2z^{\prime}=\|u^{\prime}+v^{\prime}\|^{2}$ tends to $4\|u_{0}\|^{2}$ and $w^{\prime}-2z^{\prime}=\|u^{\prime}-v^{\prime}\|^{2}$ tends to 0. Therefore, $(u^{\prime},v^{\prime})$ tends to $\left\{u=v,\|u\|=\|u_{0}\|\right\}$. Note, the map $T$ is invariant under rotation. Therefore, with proper rotation, we can select $(u^{\prime},v^{\prime})$ such that, as $(z^{\prime},w^{\prime})$ tends to $(\|u_{0}\|^{2},2\|u_{0}\|^{2})$, it tends to $\left\{u=v,u=u_{0}\right\}=(u_{0},u_{0})$. Thus, such $(u^{\prime},v^{\prime})$ lies in $W$. This gives the claim.
Finally, by Proposition 26, there exist $\theta^{\prime},\theta^{\prime\prime}\in W$ such that $\mathrm{GD}_{\eta}^{N}(\theta^{\prime})$ converges to $p^{+}(\theta^{\prime})$ and $\mathrm{GD}_{\eta}^{N}(\theta^{\prime\prime})$ converges to $p^{-}(\theta^{\prime})$. The case of $y<0$ can be proved analogously using Proposition 26. This completes the proof. ∎
Proof of Theorem 4.
The results directly come from Proposition 8 and Proposition 22. ∎